### v2

# FractalForensics: Proactive Deepfake Detection and Localization via Fractal Watermarks

Abstract.
Proactive Deepfake detection via robust watermarks has seen interest ever since passive Deepfake detectors encountered challenges in identifying high-quality synthetic images. However, while demonstrating reasonable detection performance, they lack localization functionality and explainability in detection results. Additionally, the unstable robustness of watermarks can significantly affect the detection performance. In this study, we propose novel fractal watermarks for proactive Deepfake detection and localization, namely FractalForensics. Benefiting from the characteristics of fractals, we devise a parameter-driven watermark generation pipeline that derives fractal-based watermarks and performs one-way encryption of the selected parameters. Subsequently, we propose a semi-fragile watermarking framework for watermark embedding and recovery, trained to be robust against benign image processing operations and fragile when facing Deepfake manipulations in a black-box setting. Moreover, we introduce an entry-to-patch strategy that implicitly embeds the watermark matrix entries into image patches at corresponding positions, achieving localization of Deepfake manipulations. Extensive experiments demonstrate satisfactory robustness and fragility of our approach against common image processing operations and Deepfake manipulations, outperforming state-of-the-art semi-fragile watermarking algorithms and passive detectors for Deepfake detection. Furthermore, by highlighting the areas manipulated, our method provides explainability for the proactive Deepfake detection results [^1] Code will be available at https://github.com/wangty1/fractalforensics..
Deepfake Detection, Deepfake Manipulation Localization, Semi-Fragile Watermarks, Digital Forensics
1. Introduction
The rapid development of Deepfake image manipulation techniques has introduced significant risks to individuals’ cyber privacy and safety, despite offering benefits to society in areas such as entertainment and education (Wang et al., 2024d). Since passive Deepfake detection methods (Rössler et al., 2019; Luo et al., 2023; Shiohara and Yamasaki, 2022; Kong et al., 2024; Dong et al., 2023; Cao et al., 2022; Luo et al., 2024) often face performance bottlenecks when encountering previously unseen generative algorithms, recent studies have begun to explore proactive defenses that intervene prior to image tampering.
Figure 1. Pipeline of watermark generation and encryption based on user-chosen parameters. Given a standard Hilbert curve, parameters $r$, $m$, and $o$ determine rotation, mirroring, and order modification variations when deriving the raw watermark, while $x_{0}$, $a$, $k$, and $d$ determine how the watermark is securely encrypted in the chaotic system.
Proactive defense using watermarks can be broadly categorized into robust and semi-fragile methods. Although both aim to insert predefined watermarks for image protection and ensure robustness under common image processing operations, the former is designed to maintain robustness even under malicious Deepfake manipulations (Wang et al., 2024a, b), while the latter intentionally allows the watermarks to degrade (Yang et al., 2021; Neekhara et al., 2024). However, existing methods face several key limitations. First, to consistently retain watermark integrity, robust watermarking frameworks usually require training with Deepfake manipulations. Therefore, the robustness tends to significantly degrade when encountering unseen Deepfake manipulation categories [^2] E.g., LampMark (Wang et al., 2024b) has an average of 96.46% watermark recovery rate against face swapping algorithms (trained with) but an average of 73.61% against face reenactment algorithms (unseen). , undermining the reliability of detection results. On the other hand, while allowing fragility when recovering watermarks from Deepfake-manipulated images, the semi-fragile approaches are usually vulnerable to common image processing operations that imitate real-life scenarios. Furthermore, to support forensic investigations in practical Deepfake-related criminal cases, beyond Deepfake detection, the current methods lack explainable localizations of the Deepfake-manipulated area within suspect images. Lastly, since most semi-fragile methods mainly focus on the watermarking frameworks, the watermark confidentiality and ease of use have been barely discussed.
To address the aforementioned issues, we propose FractalForensics, a semi-fragile watermarking approach for proactive Deepfake detection and localization. First, we design a parameter-driven watermark generation pipeline (Figure 1). Specifically, to achieve the localization functionality, we use parameterized fractal curves (Allezaud, 1985) to generate watermarks with spatially meaningful structures and minimal storage overhead. Unlike randomly defined watermark patterns, we devise fractal-based watermarks that are generated following deterministic rules and a small set of parameters, allowing on-demand reconstruction without storing ground-truth layouts. Meanwhile, the recursive and space-filling nature of fractals ensures that watermarks span across most regions of images, which is essential for localizing Deepfake-manipulated areas. On top of the structural backbone, we apply variations to diversify fractal watermark shapes, and to guarantee watermark confidentiality, we further utilize a chaotic encryption system (Matthews, 1989) to perform one-way encryption on fractal watermarks, ensuring unpredictability and irreversibility. In general, the entire pipeline is parameter-driven and confidential by design, eliminating the need to store raw watermark data.
Then, we construct an end-to-end watermarking framework for embedding and recovery (Figure 2). To preserve spatial correspondence and enable localization, each fractal watermark is represented as a 2-dimensional matrix with 4-bit entries, each encoded as a 4-channel binary tensor. We introduce an entry-to-patch watermark embedding strategy that maps each watermark entry to the image patch at the corresponding position. In particular, for the input image, an image feature extraction module obtains critical image information and a watermark diffusion module expands and refines the input watermark, where outputs of the two modules are passed to the watermark fusion module for watermarked image reconstruction. Thereafter, we construct a decoder for watermark recovery with expected robustness and fragility. In the end, proactive Deepfake detection and localization are accomplished based on the recovered watermarks. The entire framework is trained in a black-box condition without Deepfake manipulations. Extensive experiments further establish the satisfactory Deepfake detection and localization ability of the proposed FractalForensics. The contributions of this work are threefold:
  • We propose FractalForensics, a novel watermarking framework via semi-fragile fractal watermarks for proactive Deepfake detection and localization in the black-box scenario.
  • Based on the characteristics of fractals, we devise a parameter-driven watermark generation pipeline and introduce an entry-to-patch watermarking strategy that implicitly ensures explainable localization of the manipulated area.
  • Experiments demonstrate promising robustness and fragility of our watermarks under common and Deepfake manipulations, outperforming state-of-the-art passive and semi-fragile approaches. Additionally, our method generalizes well to unseen datasets and Deepfake manipulations.
2. Related Work
2.1. Watermarks for Deepfake Detection
Although classic passive detectors (Rössler et al., 2019; Wang and Chow, 2023; Shiohara and Yamasaki, 2022; Wang et al., 2023; Kong et al., 2022, 2025) have achieved significant progress since the first occurrence of Deepfake, recent advancements in generative models (Wang et al., 2024c; Xu et al., 2022; Zhao et al., 2023b; Tan et al., 2024) have brought strong challenges with high-quality synthetic images. As countermoves, proactive approaches (Huang et al., 2022; Qu et al., 2024; Wang et al., 2025; Wu et al., 2023; Wang et al., 2024a) insert invisible signals into benign images in advance of Deepfake manipulations and actively determine real or fake based on the signals’ functionality afterwards.
Figure 2. Workflow of the proposed FractalForensics. The channel-wise watermark matrix $w^{c}$ is embedded into image II following image feature extraction, watermark diffusion, and watermark fusion modules. To recover the watermark from $I_{\textrm{rec}}$, we construct a decoder. The framework is trained end-to-end. In the testing phase, Deepfake localization is implicitly achieved.
Robust watermarking frameworks (Zhu et al., 2018; Jia et al., 2021; Ma et al., 2022) are traditionally used for copyright verification and authentication, while recent work assigns semantics to the robust watermarks for Deepfake detection. In 2021, FakeTagger (Wang et al., 2021) raises the general idea of embedding robust and meaningful messages into images for proactive Deepfake detection and provenance. Wang et al. (Wang et al., 2024a) proposed identity perceptual watermarks that are robustly embedded and recovered for the detection of Deepfake face swapping and source tracing based on consistency in identity information. Later, LampMark (Wang et al., 2024b) exploits structure-sensitive characteristics and introduces landmark perceptual watermarks, performing proactive Deepfake detection for both face swapping and face reenactment. However, being trained only against the face swapping model, SimSwap (Chen et al., 2020), its robustness drops significantly when evaluated against unseen face reenactment algorithms.
Semi-fragile watermarks are designed to be robust and fragile regarding common image processing operations and Deepfake manipulations, respectively. An early work, FaceGuard (Yang et al., 2021), raises the idea of assigning fragility to watermarks when facing Deepfake manipulations. Later, Kong et al. (Kong et al., 2022) proposed an innovative dual-branch architecture and achieved state-of-the-art performance in Deepfake localization. WaterLo (Beuve et al., 2023) uniformly applies a 1-bit message to the entire benign image and achieves localization for early face swapping algorithms depending on the presence of watermarks in the facial area. Zhao et al. (Zhao et al., 2023a) inserted a watermark into the identity features of the images, and upon modifying facial identities by face swapping manipulations, the image authenticity is determined based on the absence of the watermark. Similarly, FaceSigns (Neekhara et al., 2024) embeds predefined 128-bit semi-fragile binary watermarks into images via a UNet architecture for Deepfake detection. Recently, EditGuard (Zhang et al., 2024) and OmniGuard (Zhang et al., 2025) treat this task as a joint image-bit steganography issue and embed semi-fragile watermarks that accomplish detection and localization regarding AI-generated contents (AIGC) simultaneously. Additionally, some recent studies (Sun et al., 2024, 2022; Wu et al., 2023) propose to utilize robust and semi-fragile watermarks together and achieve proactive detection based on the co-existence of the watermarks, where their semi-fragile watermarks are also worth comparison and discussion in this study.
2.2. Fractals Patterns for Structural Coverage
Fractals (Allezaud, 1985) have been widely studied for their space-filling properties to traverse all points in a 2-dimensional grid (Mandelbrot, 1982). While exhibiting strong spatial continuity, self-similarity, and locality preservation, these curves have been proven to be suitable for applications in image compression (Reza and Flickner, 1999) and texture synthesis (Woop et al., 2011). However, their potential for structured watermark design in the context of proactive Deepfake detection remains largely unexplored. In this work, leveraging the fractal patterns, we derive parameter-driven watermark matrices that are spatially coherent, enabling implicit localization besides proactive Deepfake detection.
3. Methodology
3.1. Overview
Deepfakes manipulate facial areas in full image frames. To protect the faces, we propose FractalForensics, a semi-fragile watermarking framework for proactive Deepfake detection and localization. For a clean facial image $I$ cropped from a full image frame, an encrypted fractal watermark matrix $w$ is generated via a parameter-driven pipeline (Figure 1). This pipeline incorporates user-selected parameters that define the fractal type, structural variation, and chaotic encryption keys. Subsequently, a watermarking framework (Figure 2) is established to invisibly embed $w$ into $I$, consisting of an image feature extraction module, a watermark diffusion module, and a watermark fusion module, to produce the watermarked image $I_{\textrm{rec}}$. A decoder is then constructed to recover the watermark from $I_{\textrm{rec}}$ as $w_{\textrm{rec}}$. By comparing $w_{\textrm{rec}}$ with the original $w$, we achieve proactive Deepfake detection and localization.
3.2. Watermark Generation Pipeline
Fractals are recursive geometric shapes whose structures are derived based on parameters. Rather than storing all watermarks like in previous studies, we propose a watermark generation pipeline that produces encrypted watermarks depending on the selected parameters. In this study, we use the Hilbert curve (Hilbert, 1891) as the selected fractal shape for demonstration. The Hilbert curve of order $n$ defines a continuous path that maps each integer index $i\in[0,2^{2n}-1]$ to a unique coordinate in a $2^{n}\times 2^{n}$ grid. The mapping follows a recursive and self-similar rule that fills the space while preserving spatial locality. Relying on the traversal order defined by the standard Hilbert curve, we construct a 2-dimensional watermark matrix as a representation of the curve.
To ensure diversity in the selected fractal shapes, we apply shape variations toward the standard Hilbert curve by rotations $r\in$ {0: 0°, 1: 90°, 2: 180°, 3: 270°}, mirroring $m\in$ {0: None, 1: Top, 2: Bottom, 3: Left, 4: Right, 5: Top Left, 6: Top Right, 7: Bottom Left, 8: Bottom Right}, and order modifications $o\in$ {0: None, 1: Reverse, 2: Zigzag, 3: Cross Flip}, leading to 144 different unique shape variations based on the combination of parameters. For each curve, there is a start and an end point that record the order and direction. To securely protect the curve shape along with the order and direction, we utilize a chaotic encryption (Matthews, 1989) system to perform one-way encryption based on the selected parameters. In specific, the logistic map is defined by
(1) $x_{i}=ax_{i-1}(1-x_{i-1}),$
where $x_{0}\in[0.1,0.9]^{\mathbb{R}}$ and $a\in[3.7,4.0)^{\mathbb{R}}$ are parameters that determine the chaotic map. To ensure the unpredictability and irreversibility of encrypted watermarks, we let parameters $k\in[100,1000]^{\mathbb{Z}}$ and $d\in[2,20]^{\mathbb{Z}}$ determine the starting value $x_{k}$ of the chaotic sequence and the index of the selected digit of each value. We use the $2^{2n}$ consecutive values at the $d$-th digit starting from $x_{k}$ as the encrypted watermark matrix of the raw entries following the curve shape from the start to the end point. Due to the chaotic characteristic of the encryption system, unless the confidential parameters are known, attempting to recover the original watermark curve by brute force is worse than NP-Hard (van Leeuwen, 1991). For watermarking purposes, since there is more tolerance for binary entries when encoding and decoding, the encrypted watermark entries are transformed into binary formats with 4 bits for each entry.
3.3. Watermarking Framework
We construct a watermarking framework with deep neural networks, aiming to invisibly embed watermark $w$ into image $I$ and perform recovery with robustness and fragility accordingly. To achieve the localization functionality for tampered areas, we introduce an innovative entry-to-patch strategy for watermarking. Specifically, the input image $I$ is implicitly recognized as patches of size $32\times 32$, and the $2^{n}\times 2^{n}$ patches are positionally matched to the watermark matrix $w$ of the same size. To ensure locality preservation such that the watermarks are positionally embedded in the desired locations, we place the 4-bit entries on the channel dimension so that they share the same spatial positions, resulting in the channel-wise binary watermark matrix $w^{c}$ of shape $4\times 2^{n}\times 2^{n}$.
Image Feature Extraction. We construct an image feature extraction module that contains a ConvBlock with a kernel size of $7\times 7$ followed by $N_{\textrm{img}}$ SEResBlocks with kernel sizes of $3\times 3$. In particular, a ConvBlock contains a sequentially executed CNN layer, batch normalization layer, and LeakyReLU activation. An SEResBlock utilizes the ResNet (He et al., 2016) bottleneck block and applies the squeeze-and-excitation networks (Hu et al., 2018) (SENet) for efficient channel-wise attention. For image $I$, we derive critical image features, with residually preserved information from early stages by SEResBlocks, for image reconstruction in the later module. The image features are analyzed in the pixel-aligned feature domain such that spatial dimensions of intermediate image features remain unchanged to maintain strict position alignment, providing favorable conditions for entry-to-patch watermarking with minimal information losses.
Watermark Diffusion. To enable effective and robust watermark embedding, we expand and refine the binary channel-wise watermark matrix $w^{c}\in\{0,1\}^{4\times 2^{n}\times 2^{n}}$ before fusion for image reconstruction. In specific, the watermark diffusion module first upsamples $w^{c}$ to the same spatial dimension as the input image, where each 4-bit entry is spatially expanded into the corresponding $32\times 32$ patch. Thereafter, the upsampled watermark is fed to a ConvBlock, $N_{\textrm{wtm}}$ SEResBlocks, and a ConvBlock to adaptively refine locally critical watermark features via moderate kernel sizes, ensuring the robustness and invisibility of watermark signals during embedding.
Watermark Fusion. In the watermark fusion module, we integrate the image and watermark features to produce the watermarked image $I_{\textrm{rec}}$. The two features are first concatenated along the channel dimension and sequentially passed through two ConvBlocks and $N_{\textrm{rec}}$ SEResBlocks, to accomplish a feature-level reconstruction while maintaining the entry-to-patch correlation locally via moderate kernel sizes and merely channel-wise attention operations, avoiding over-coherence between patches that are far apart since this can make the watermark overly robust against Deepfake when watermark fragility is required. Dropout operations are concurrently applied to avoid possible overfitting and unexpectedly robust watermark recovery. Then, to enhance reconstruction quality and obtain the ultimate RGB image, we concatenate the reconstructed feature-level result with the original image $I$ and pass them through two consecutive ConvBlocks to generate the watermarked image $I_{\textrm{rec}}$ with three channels.
Decoder. A lightweight decoder is designed for watermark recovery. In specific, a sequence of a ConvBlock, $N_{\textrm{dec}}$ SEResBlocks, and two ConvBlocks that gradually reduce spatial resolution and distill deterministic watermark information are applied to the watermarked image $I_{\textrm{rec}}$. With moderate kernel sizes, the decoder preserves positional correspondence regarding the entry-to-patch watermarking strategy, and stride-based downsampling is employed instead of pooling operations. A final Sigmoid activation followed by rounding is applied to map the output to binary values. In this end-to-end pipeline, the decoder is jointly trained with the former modules for watermark embedding, which benefits the thorough learning of the entry-to-patch scheme and balances the performance regarding visual quality and watermark robustness.
Discriminator. In this work, a discriminator is trained from scratch along with the main FractalForensics framework. Specifically, a ConvBlock and $N_{\textrm{rec}}$ SEResBlocks are applied simply followed by a binary classifier to determine whether it is watermarked or not. The discriminator is used to train the main framework adversarially.
Table 1. Visual quality evaluation of the watermarked images. Information includes model name, resolution (Res.), watermark size (Num. Bits), PSNR, SSIM, and LPIPS.
[…]
WaterLo (Beuve et al., 2023) 512 1 46.73 0.993 0.001
SepMark (Wu et al., 2023) 128 30 31.51 0.965 0.014
256 128 32.08 0.945 0.046
FaceSigns (Neekhara et al., 2024) 256 128 30.60 0.945 0.033
EditGuard (Zhang et al., 2024) 512 64 36.93 0.841 0.021
FractalForensics 256 256 35.926 0.974 0.013
512 1024 38.129 0.981 0.014
3.4. Loss Functions
Table 2. Watermark recovery rate against common image operations on CelebA-HQ.
Models WaterLo (Beuve et al., 2023) SepMark (Wu et al., 2023) FaceSigns (Neekhara et al., 2024) EditGuard (Zhang et al., 2024) FractalForensics (patch) FractalForensics (bit)
Resolution 512 128 256 256 256 256 512 256 512
Identity 99.99% 99.99% 99.99% 99.84% 99.98% 99.99% 99.99% 99.99% 99.99%
Jpeg 61.72% 99.93% 99.93% 96.68% 80.01% 99.97% 99.84% 99.99% 99.95%
Gaussian Noise 64.20% 54.47% 49.38% 52.08% 58.39% 99.76% 99.04% 99.93% 99.68%
Gaussian Blur 61.76% 99.99% 99.99% 98.52% 0.23% 99.57% 99.96% 99.88% 99.99%
Median Blur 89.25% 99.99% 99.99% 98.44% 3.86% 99.99% 99.97% 99.96% 99.99%
Resize 61.70% 99.99% 99.99% 96.73% 3.54% 98.03% 99.56% 99.47% 99.86%
[…]
In this study, we focus on two main objectives: visual quality and watermark recovery performance. We first assign a pixel-level Mean Squared Error (MSE) loss on $I$ and $I_{\textrm{rec}}$ by
(2) $\mathcal{L_{\textrm{MSE}}}=\lVert I-I_{\textrm{rec}}\rVert^{2}_{2}.$
Meanwhile, to make sure that $I_{\textrm{rec}}$ is perceptually identical to $I$ for human eyes, we adopt the Learned Perceptual Image Patch Similarity (LPIPS) loss following
(3) $\mathcal{L_{\textrm{LPIPS}}}=\Sigma_{l}\lVert\phi_{l}(I)-\phi_{l}(I_{\textrm{rec}})\rVert_{2}^{2},$
where $\phi_{l}$ represents the feature activation from layer $l$ of a pre-trained AlexNet (Krizhevsky et al., 2012). Meanwhile, the discriminator $D$ is trained to distinguish $I$ and $I_{\textrm{rec}}$ following
(4) $\mathcal{L}_{D}=-\mathbb{E}(\log D(I))+\mathbb{E}\log(1-D(I_{\textrm{rec}})),$
and it assists in training the watermarking framework for better visual quality following
(5) $\mathcal{L}_{\textrm{adv}}=-\mathbb{E}(\log D(I_{\textrm{rec}})).$
We train the decoder for watermark robustness following
(6) $\mathcal{L}_{\textrm{dec}}=\lVert m_{\textrm{rec}}-m\rVert.$
In summary, the total loss is computed following
(7) $\mathcal{L}=\lambda_{\textrm{MSE}}\mathcal{L_{\textrm{MSE}}}+\lambda_{\textrm{LPIPS}}\mathcal{L_{\textrm{LPIPS}}}+\lambda_{\textrm{adv}}\mathcal{L}_{\textrm{adv}}+\lambda_{\textrm{dec}}\mathcal{L}_{\textrm{dec}}.$
4. Experiments
4.1. Implementation Details
We used the popular facial image datasets, CelebA-HQ (Karras et al., 2018) and LFW (Huang et al., 2012), for experiments. Specifically, we followed the official split of CelebA-HQ with 30,000 facial images for training, validation, and testing, and we used LFW with 5,749 facial identities for evaluating the generalizability. We trained two versions of FractalForensics for the 256 and 512 resolutions, with $N_{\textrm{img}}$ being 5 and 6, $N_{\textrm{wtm}}$ being 5 and 4, $N_{\textrm{rec}}$ being 5 and 4, and $N_{\textrm{dec}}$ being 4 and 3, respectively. For the loss weights, we set $\lambda_{\textrm{MSE}}=1$, $\lambda_{\textrm{LPIPS}}=0.5$, $\lambda_{\textrm{adv}}=0.01$, and $\lambda_{\textrm{dec}}=12$. During training, we used Identity, Jpeg, and Gaussian Noise as common image processing operations to boost watermark robustness. Deepfake manipulations are completely omitted during training under the black-box setting. In the testing phase, the semi-fragile watermarking algorithms are evaluated against common image processing operations including Identity, Jpeg, Gaussian Noise, Gaussian Blur, Median Blur, and Resize and Deepfake manipulations including SimSwap (Chen et al., 2020), InfoSwap (Gao et al., 2021), UniFace (Xu et al., 2022), E4S (Liu et al., 2023), and DiffSwap (Zhao et al., 2023b) for face swapping and StarGAN (Choi et al., 2020), StyleMask (Bounareli et al., 2023b), and HyperReenact (Bounareli et al., 2023a) for face reenactment. We trained the main watermarking framework at a learning rate of $0.002$ and trained the discriminator at $0.0004$. All experiments are conducted on 8 Tesla A100 GPUs.
4.2. Watermarking Performance on CelebA-HQ
In this section, we evaluate the watermarking performance on CelebA-HQ (Karras et al., 2018), with respect to the visual quality of the watermarked images and the watermark robustness and fragility with respect to common image operations and Deepfake manipulations.
Visual Quality. We conducted quantitative evaluations of the visual quality after embedding watermarks. Specifically, we compared our proposed FractalForensics to state-of-the-art semi-fragile watermarks for proactive defense against Deepfake and AIGC. As Table 1 reports, our method achieves the second-best performance with the most difficult bit-per-pixel rate, while the best WaterLo (Beuve et al., 2023) has watermarks with only 1 bit. In general, while all methods demonstrate reasonable LPIPS values, SepMark [^3] SepMark consists of a robust and a semi-fragile watermark, where we adopted the semi-fragile watermark for comparing robustness and fragility in this study. (Wu et al., 2023) and FaceSigns (Neekhara et al., 2024) both have PSNR values below 35, and EditGuard (Zhang et al., 2024) derives an unsatisfactory SSIM value below 0.85. As a result, our approach consistently maintains promising visual qualities at both resolutions for all evaluation metrics on visual quality.
Table 3. Watermark recovery rate against Deepfake image manipulations on CelebA-HQ.
Models WaterLo (Beuve et al., 2023) SepMark (Wu et al., 2023) FaceSigns (Neekhara et al., 2024) EditGuard (Zhang et al., 2024) FractalForensics
Resolution 512 128 256 256 256 256 512
SimSwap (Chen et al., 2020) 61.72% 98.97% 69.49% 50.21% 49.23% 78.19% 60.79%
InfoSwap (Gao et al., 2021) 62.10% 99.94% 98.38% 50.46% 37.04% 75.12% 72.55%
UniFace (Xu et al., 2022) 61.74% 50.11% 49.86% 49.95% 49.18% 64.66% 24.11%
E4S (Liu et al., 2023) 62.34% 99.78% 95.33% 50.19% 50.05% 69.19% 66.81%
DiffSwap (Zhao et al., 2023b) 62.08% 50.23% 49.79% 51.79% 50.73% 81.31% 59.91%
StarGAN (Choi et al., 2020) 60.89% 49.93% 50.07% 50.17% 49.90% 9.92% 10.02%
StyleMask (Bounareli et al., 2023b) 66.27% 49.86% 50.10% 49.98% 50.14% 10.00% 9.96%
HyperReenact (Bounareli et al., 2023a) 61.88% 50.11% 50.06% 50.13% 49.87% 9.98% 10.02%
[…]
Table 4. Deepfake detection performance in AUC scores on CelebA-HQ.
Model Xception (Rössler et al., 2019) SBIs (Shiohara and Yamasaki, 2022) RECCE (Cao et al., 2022) CADDM (Dong et al., 2023) WaterLo (Beuve et al., 2023) SepMark (Wu et al., 2023) FaceSigns (Neekhara et al., 2024) EditGuard (Zhang et al., 2024) FractalForensics
SimSwap (Chen et al., 2020) 71.15% 88.94% 69.01% 87.66% 69.77% 99.97% 98.87% 49.33% 99.99%
InfoSwap (Gao et al., 2021) 65.50% 80.50% 52.13% 61.39% 51.30% 64.83% 98.87% 50.36% 99.99%
UniFace (Xu et al., 2022) 70.34% 79.41% 67.35% 82.73% 69.69% 99.99% 98.91% 46.71% 99.99%
E4S (Liu et al., 2023) 53.70% 61.05% 47.19% 73.13% 63.99% 73.12% 98.87% 56.37% 99.99%
DiffSwap (Zhao et al., 2023b) 53.62% 56.15% 63.60% 73.33% 52.03% 99.99% 98.83% 48.56% 99.99%
StarGAN (Choi et al., 2020) 49.30% 65.86% 41.55% 44.34% 68.56% 99.99% 98.87% 47.12% 99.99%
StyleMask (Bounareli et al., 2023b) 40.23% 48.45% 23.87% 39.73% 50.12% 99.99% 98.87% 47.03% 99.99%
HyperReenact (Bounareli et al., 2023a) 76.27% 53.35% 78.23% 42.87% 64.86% 99.99% 98.88% 47.03% 99.99%
Average 60.01% 66.71% 55.37% 63.15% 61.29% 92.23% 98.87% 49.06% 99.99%
Watermark Recovery Rate. We evaluated the watermark recovery rate of our approach against common image processing operations and Deepfake manipulations, compared to contrastive semi-fragile watermarking frameworks. As demonstrated in Table 2, while all algorithms stay robust as expected when applied with no operation (Identity), they each suffer a drastic reduction in watermark recovery rates when facing some image processing operations. In general, Gaussian Noise has the most serious effect, causing all watermark recovery rates to fall below 65% for the comparative algorithms. Meanwhile, the two models that have demonstrated relatively better visual quality in Table 1, WaterLo (Beuve et al., 2023) and EditGuard (Zhang et al., 2024), also fail to maintain robustness regarding Jpeg, Gaussian Blur, Median Blur, and Resize operations. For our proposed FractalForensics, since a 4-bit binary value represents each watermark entry, we computed the patch-wise watermark recovery rate as our official performance and used that for Deepfake detection and localization in later sections, which is harder than the bit-wise metric in preserving robustness. As a result, our method achieves the best performance on average with 99.55% and 99.73% at the 256 and 512 resolutions. In the last two columns, we also presented the bit-wise watermark recovery rates of FractalForensics as references, where reasonably higher values than the patch-wise ones can be observed as expected since the patch-wise watermark recovery rate is computed following a more strict rule.
When tested against Deepfake manipulations, the watermark fragility is expected to be observed. In other words, there is expected to be an obvious gap between the watermark recovery rates for robustness and fragility, such that real and fake samples are distinguishable accordingly. For this experiment, we adopted five face swapping (Chen et al., 2020; Gao et al., 2021; Xu et al., 2022; Liu et al., 2023; Zhao et al., 2023b) and three face reenactment (Choi et al., 2020; Bounareli et al., 2023b, a) manipulations for evaluation. It can be observed from Table 3 that the fragility has been achieved by WaterLo (Beuve et al., 2023), FaceSigns (Neekhara et al., 2024), EditGuard (Zhang et al., 2024), and our proposed FractalForensics, while SepMark (Wu et al., 2023) demonstrates robust watermark recovery ability from images generated by SimSwap (Chen et al., 2020), InfoSwap (Gao et al., 2021), and E4S (Liu et al., 2023). Meanwhile, FractalForensics is observed to demonstrate promising fragility at both resolutions with average watermark recovery rates of 50.05% and 39.27%.
With a closer look at Table 2 and Table 3, since there exist watermark recovery rates against common image operations that are lower than those against Deepfake manipulations for the comparative models (e.g., Sepmark has 99.94% for InfoSwap but 54.47% for Gaussian Noise), it becomes impossible to distinguish real and fake based on robustness and fragility. As a result, although our FractalForensics demonstrates slightly higher watermark recovery rates than the other semi-fragile watermarks when tested against face swapping algorithms, there remains a clear gap in the recovery rates regarding robustness and fragility. Additionally, since our watermark matrices are embedded following an entry-to-patch strategy for localizing Deepfake manipulations, only the edited patches are expected to lose the embedded watermarks. Therefore, the critically lower watermark recovery rates for face reenactment than those for face swapping are reasonable and consistent with the localization results in Section 4.4.
4.3. Deepfake Detection
Considering the robustness and fragility of watermarks regarding real and Deepfake images, semi-fragile watermarks act as a natural proactive Deepfake detector. In particular, an ideal case is expected to showcase watermark recovery rates of Deepfake images all lower than those of real images, demonstrating a clear threshold between real and fake samples. In this study, following this scheme, we evaluated the Deepfake detection performance of semi-fragile watermarking frameworks. Since no clear threshold has been provided for the comparative models, we adopted the AUC scores for evaluation. In addition, we included four state-of-the-art passive Deepfake detectors (Rössler et al., 2019; Shiohara and Yamasaki, 2022; Cao et al., 2022; Dong et al., 2023) that have shown reasonable performance on lab-controlled datasets for comparison. In this experiment, the real samples are derived by applying common image operations as listed in Table 2 and the fake ones are produced by performing Deepfake manipulations on the real images in CelebA-HQ.
Figure 3. Effect of image operations on the watermarked image with localization demonstrated by the red overlay.
As Table 4 lists, although demonstrating some acceptable detection results (e.g., SBIs (Shiohara and Yamasaki, 2022) and CADDM (Dong et al., 2023) on SimSwap), the passive detectors generally exhibit poor generalizability toward most Deepfake manipulations in the experiments, leading to average AUC scores below 70%. On the other hand, while most semi-fragile algorithms have exhibited desired watermark fragility against most Deepfake manipulations, the proactive Deepfake detection performance depends on their robustness against common image operations. As a result, since a clear gap between watermark recovery rates for real and fake can be observed for FractalForensics, we promisingly achieved satisfactory Deepfake detection results with an average of 99.99% AUC score over all participating Deepfake manipulations. As for the comparative ones, due to larger watermark recovery rates regarding fragility than those for robustness, WaterLo (Beuve et al., 2023) and EditGuard (Zhang et al., 2024) fail to achieve Deepfake detection with poor AUC scores. SepMark (Wu et al., 2023), although shows satisfactory performance with AUC scores of 99.99% for some Deepfake models, the surprisingly high watermark recovery rates in Table 3 result in much lower AUC scores for InfoSwap (Gao et al., 2021) and E4S (Liu et al., 2023), leading to an average of 92.23% in Table 4. Meanwhile, although the watermark recovery rates are close for robustness and fragility, there exists a small gap that benefits FaceSigns (Neekhara et al., 2024) in deriving reasonable AUC scores against all Deepfake manipulations. However, its fragility against Gaussian Noise remains a potential risk.
Figure 4. Effect of Deepfake algorithms on the watermarked image with localization demonstrated by the red overlay.
4.4. Deepfake Manipulation Localization
Localization Against Common Image Operations. We first visualized the localization results of our approach when applying common image operations to watermarked images. Since the watermarks are expected to demonstrate robustness when facing common image operations that imitate real-life cases, it is considered a good performance if localized areas barely appear. As Figure 3 demonstrates, patches that lose watermarks are highlighted with red overlays. In particular, the exhibition is convincingly consistent with the statistics reported in Table 2, such that FractalForensics is observed to be more sensitive to the Resize operation. In summary, the behavior of the watermarks matches the expectation.
Localization Against Deepfake Manipulations. We then evaluated the localization results against Deepfake manipulations. Since Deepfake images are produced upon feature fusion via deep neural networks, there is no ground-truth for localizing Deepfake manipulations. Therefore, FractalForensics is designed to embed watermarks with an entry-to-patch strategy that implicitly enables the localization functionality. As a result, as visualized in Figure 4, the first and second rows display the clean Deepfake results via different manipulation algorithms and their corresponding localization results regarding the sample pair of images, respectively. In the third row, we recorded the number of 4-bit watermark entries that were lost for each patch position on the entire testing dataset and plotted the cumulative localization results after normalization. Given the fact that CelebA-HQ contains facial images that are well-aligned with faces mostly located in the center of images, it can be concluded that, since face swapping manipulations mostly modify detailed facial organs and cues, the localized manipulated areas are mainly in the center of the images. It is also noticed that the localized areas for DiffSwap (Zhao et al., 2023b) are mainly in the bottom half of the face. This is because DiffSwap does not heavily edit the forehead and eyes but the nose and mouth. On the other hand, face reenactment algorithms edit not only facial expressions that are located within the face area but also the head poses that can heavily affect the background area, leading to localized areas that largely span entire images. Additionally, by taking a look at Table 3 together with Figure 4, consistencies can be observed between localization visualizations and watermark recovery rates.
Figure 5. Qualitative comparison of localization performance between watermarking frameworks. The localized fake areas are highlighted in purple, white, and red by WaterLo, EditGuard, and our proposed FractalForensics, respectively.
Table 5. Cross-dataset evaluation on LFW for visual quality and watermark recovery rate against common image operations.
[…]
WaterLo (Beuve et al., 2023) 99.89% 61.84% 64.53% 72.96% 99.59% 61.87% 76.78% 46.807 0.991 0.001
SepMark (Wu et al., 2023) 99.99% 99.78% 88.05% 99.99% 99.99% 99.99% 97.97% 30.825 0.947 0.021
FaceSigns (Neekhara et al., 2024) 99.98% 99.24% 66.61% 99.95% 99.91% 99.43% 94.19% 32.251 0.955 0.026
EditGuard (Zhang et al., 2024) 99.95% 78.23% 58.54% 0.10% 0.81% 1.93% 39.92% 37.880 0.749 0.031
FractalForensics 99.95% 99.53% 99.35% 98.97% 99.50% 97.16% 99.08% 35.741 0.970 0.019
Table 6. Cross-dataset evaluation on LFW for watermark recovery rate against Deepfake manipulations.
[…]
WaterLo (Beuve et al., 2023) 62.07% 45.63% 61.84% 62.42% 62.07% 62.14% 65.93% 62.81% 60.74%
SepMark (Wu et al., 2023) 69.57% 98.14% 52.39% 99.87% 50.16% 50.00% 49.68% 50.02% 64.31%
FaceSigns (Neekhara et al., 2024) 50.08% 50.31% 49.89% 50.03% 50.71% 50.03% 49.92% 50.28% 56.97%
EditGuard (Zhang et al., 2024) 49.35% 42.48% 49.16% 50.35% 50.49% 50.05% 49.94% 49.99% 65.30%
FractalForensics 74.87% 68.57% 51.49% 82.60% 91.48% 9.97% 10.00% 10.08% 49.88%
Furthermore, as WaterLo (Beuve et al., 2023) and EditGuard (Zhang et al., 2024) are also designed to have localization functionality against Deepfake and AIGC, we conducted comparative experiments for the models. The localization results are shown in Figure 5. Specifically, WaterLo defines purple overlays as fake areas and uses green for real, while EditGuard highlights the fake and real areas with white and black colors, respectively. However, although successfully highlighting the entire image as fake for all Deepfake-manipulated images, they both lack detailed information that highlights specific manipulated areas. On the other hand, since they each are vulnerable to some common image operations, the localization results become relatively unconvincing. In contrast, Figures 34, and 5 show the reliability of FractalForensics in localizing Deepfake manipulated areas and provide explainable evidence to support Deepfake detection.
Watermark Sensitivity Against Cropping. In this study, since the watermarking pipeline follows an entry-to-patch strategy, as a side contribution, our approach also exhibits sensitivity against malicious cropping. To explore the watermark robustness in this case, we conducted experiments to randomly crop adjacent image patches and verify whether watermarks in the cropped areas were lost and those in the untampered areas remained. As demonstrated in Figure 6, we attempted cropping sizes from $1\times 1$ to $5\times 5$ patches, and we recorded the bit-wise and patch-wise watermark recovery rates, along with the correctness regarding desired robustness and fragility. As a result, the bit-wise and patch-wise watermark recovery rates exhibit declines that generally match with the number of cropped patches for both CelebA-HQ and LFW, while the overall correctness stays above 95% even when 25 out of 64 image patches are cropped within an image.
4.5. Cross-Dataset Evaluation on LFW
In this section, to evaluate the cross-dataset performance of our proposed FractalForensics, we conducted experiments on LFW (Huang et al., 2012). Specifically, experiments follow the same settings as in Table 2 and Table 3 to compute the watermark recovery rates against common image operations and Deepfake manipulations. For simplicity, the 256 resolution versions for SepMark (Wu et al., 2023) and our FractalForensics are utilized in the comparison.
As Table 5 reports, the comparative watermarking frameworks generally maintain consistent performance as for CelebA-HQ (Karras et al., 2018), demonstrating obvious vulnerability when facing Gaussian Noise, with no watermark recovery rate above 90%. In contrast, although still slightly affected by the Resize operation, our proposed FractalForensics generally preserves robustness against all common image operations at an average of 99.08% with good visual quality. Additionally, while showing outstanding visual quality, WaterLo fails to achieve the expected watermark robustness. Besides the unsatisfactory robustness and a promising PSNR value, EditGuard exhibits an unexpectedly low SSIM value, which may be caused by mismatched image color or tone.
Similar outcomes can be seen in Table 6 when tested against Deepfake manipulations, such that the comparative models mostly demonstrate fragility while SepMark (Wu et al., 2023) is surprisingly robust against InfoSwap (Gao et al., 2021) and E4S (Liu et al., 2023). Consequently, the proposed FractalForensics obtains an optimal average watermark recovery rate of 49.88% in this comparison. Furthermore, when taking a look at Table 5 and Table 6 together, a clear gap for watermark robustness and fragility regarding common image operations and Deepfake manipulations can still be found for our proposed FractalForensics. In conclusion, our approach performs satisfactorily even under the cross-dataset setting, outperforming the state-of-the-art semi-fragile watermarking frameworks.
Figure 6. Watermark robustness and localization performance against cropping on CelebA-HQ and LFW.
5. Conclusion
In this work, we propose FractalForensics, a novel proactive Deepfake detection method based on semi-fragile fractal watermarks. Relying on the characteristics of fractal shapes, we first design a parameter-driven pipeline for watermark generation, allowing user-selected parameters to determine the variations toward the fractal shape and the watermark encryption rule via the chaotic encryption system. Upon placing the 4-bit watermark matrix entries at the channel dimension, we introduce an innovative entry-to-patch strategy for position-aware watermark embedding and recovery, which implicitly helps accomplish explainable localizations of malicious Deepfakes. In conclusion, FractalForensics achieves excellent robustness and fragility for the watermark recovery rates, leading to promising proactive Deepfake detection and localization performance, outperforming state-of-the-art passive and semi-fragile approaches. Future work may continue to exploit the self-similarity of fractal shapes, where an automatic self-check can be applied to the recovered watermark without knowing the parameters. If capacity allows, an auxiliary robust watermark may also be integrated for additional source-tracing functionality. Moreover, since real-life cases are complex and contain not only facial parts, further developments are still required for actual deployment [^4] This work has been accepted to ACM Multimedia 2025 as an oral paper. .
Acknowledgements.
This research is supported by the Ministry of Education, Singapore, under its MOE AcRF TIER 3 Grant (MOE-MOET32022-0001) and by VoiceCode Technology Pte. Ltd.
Appendix A Extra Experimental Results
In this section, we report extra experimental results in addition to those discussed in the main content.
A.1. Visual Quality.
In Figure 7, with regard to the reference faces, we applied each Deepfake manipulation to the clean and watermarked victim images in every two columns. In other words, we evaluated whether the embedded watermarks heavily affect the Deepfake manipulation results. It can be observed that, although trivial differences appear [^5] DiffSwap (Zhao et al., 2023b) contains randomness upon every run, leading to differences even without watermarks. , the Deepfake manipulations are generally unaffected, allowing potential benign utilization.
Figure 8. Watermarked images produced by each semi-fragile watermarking framework.
In Figure 8, we visualized some watermarked images produced by each semi-fragile watermarking framework. It can be observed that the results are generally consistent with the quantitative visual quality evaluation in the main content. In particular, WaterLo (Beuve et al., 2023) and EditGuard (Zhang et al., 2024) demonstrate outstanding PSNR values, with noise artifacts barely appearing in the images. On the other hand, obvious artifacts and distortions caused by the watermarks are displayed in rows 3 and 4. Additionally, mismatches in color and tone can be found in the row of EditGuard (column 7 being the most obvious), which matches the low SSIM values as listed in the main content.
A.2. Cross-Dataset Evaluation on LFW
Table 7. Extra cross-dataset evaluation of watermarked recovery rate against common image operations on LFW.
[…]
SepMark (Wu et al., 2023) 128 99.34% 96.69% 86.71% 99.61% 99.91% 99.74% 97.00%
FractalForensics 512 99.99% 99.73% 98.40% 99.97% 99.99% 99.78% 99.64%
Table 8. Extra cross-dataset evaluation of watermarked recovery rate against Deepfake manipulations on LFW.
[…]
SepMark (Wu et al., 2023) 128 97.72% 99.10% 53.29% 99.71% 50.19% 49.90% 49.68% 50.04% 68.69%
FractalForensics 512 61.32% 63.96% 27.62% 66.83% 89.97% 9.98% 10.00% 9.99% 42.46%
In the main content, only the 256 resolution setting is adopted for SepMark (Wu et al., 2023) and FractalForensics for simplicity. In this section, we supplement the 128 resolution for SepMark and 512 resolution for FractalForensics. As listed in Table 7 and Table 8, our proposed FractalForensics consistently performs with expected robustness against common image processing operations and fragility against Deepfake manipulations. Meanwhile, SepMark (Wu et al., 2023) generally behaves well at the 128 resolution except for a watermark recovery rate below 90% for Gaussian Noise. Moreover, as a clear gap can be found between the robustness and fragility of our approach at the 512 resolution, the performance of SepMark when facing SimSwap (Chen et al., 2020), InfoSwap (Gao et al., 2021), and E4S (Liu et al., 2023) is unsatisfactory with watermark recovery rates as high as above 97%. It is also worth noting that, for our proposed FractalForensics, since DiffSwap (Zhao et al., 2023b) modifies relatively fewer patches compared to other synthetic models and facial areas in LFW (Huang et al., 2012) occupy fewer patches than those of CelebA-HQ (Karras et al., 2018), greater watermark recovery rates can be observed when tested against DiffSwap. However, this is reasonable and does not affect the gap between robustness and fragility.
Figure 9. Cumulative localization visualizations of our FractalForensics against Deepfake manipulations on CelebA-HQ at 512 resolution and on LFW at 256 and 512 resolutions.
Figure 10. Localization performance against common image operations. Fragility can be found for WaterLo (Beuve et al., 2023) and EditGuard (Zhang et al., 2024) for most operations.
A.3. Deepfake Localization
In addition to the statistical watermark recovery rates for each experimental setting and the localization visualization for the 256 resolution on CelebA-HQ, in this section, we visualized the cumulative localization results of FractalForensics on CelebA-HQ (Karras et al., 2018) at the 256 resolution and on LFW (Huang et al., 2012) at 256 and 512 resolutions. In Figure 9, localized areas with highly mismatched watermarks are centered in the squares for face swapping models, which matches the functionality of face swapping algorithms. This makes sense since both CelebA-HQ and LFW contain facial images that have the faces aligned and mostly placed at the center while faces in LFW occupy smaller areas. Meanwhile, depending on the specific algorithm and model workflow, the amount of tampered patches varies for each swapping algorithm. As for the face reenactment models, since they edit head poses that affect the background area, most algorithms tend to smooth and blur the entire background area, leading to losing watermarks for those patches. In general, our proposed FractalForensics achieves promising and explainable localization for proactive Deepfake detection, regardless of datasets and image resolutions.
Furthermore, we conducted detailed analyses regarding the localization results of WaterLo (Beuve et al., 2023) and EditGuard (Zhang et al., 2024) for comparisons with our FractalForensics. In Figure 10, WaterLo correctly identifies the image with no operation (Identity in column 1) as unmodified with a full overlay in green. However, it is highly affected by common operations and fails to maintain the desired watermark when facing Jpeg, Gaussian Noise, Gaussian Blur, and Resize. Similarly, EditGuard correctly localizes nothing in the first column, with a full dark localization result, but unexpectedly marks all common image operations as malicious edits with large white areas.
Additionally, the localization performance of EditGuard (Zhang et al., 2024) is controlled by a threshold $\tau$ that justifies the intensity of the watermark difference between recovered and ground-truth. In Figure 11, we visualized the localization results of EditGuard on Deepfake-manipulated images by gradually increasing the threshold $\tau$ from 0.2 up to 0.8. Note that white and black represent edited and untampered regions, respectively. It can be discovered that the white areas suddenly turn into black from $\tau=0.6$ to $\tau=0.8$. As a result, although the localization result varies against different Deepfake manipulations, for EditGuard, there lacks a reasonable localization result that clearly highlights the Deepfake-manipulated region rather than annotating the entire image as fake.
Figure 11. Localization results of EditGuard as the threshold $\tau$ changes from 0.2 to 0.8.
Table 9. Watermark recovery performance against cropping areas of different sizes on CelebA-HQ.
Num. Patches Rec. Rate (bit) Rec. Rate (patch) Correctness Rem. Portion
$1\times 1$ 99.45% 98.79% 99.64% 98.44%
$2\times 2$ 97.63% 94.79% 98.95% 93.75%
$3\times 3$ 94.44% 87.89% 98.03% 85.94%
$3\times 4$ 92.51% 83.74% 97.49% 81.25%
$4\times 4$ 90.03% 78.32% 96.66% 75.00%
$4\times 5$ 87.42% 72.59% 96.14% 68.75%
$5\times 5$ 84.32% 65.76% 95.15% 60.94%
Random 95.09% 89.30% 98.19%
A.4. Watermark Performance Against Cropping
In the main content, it is proven that our proposed FractalForensics demonstrates side contributions with reasonable sensitivity regarding image cropping. In this section, in addition to the figure, in Table 9 and Table 10, we listed the detailed bit-wise watermark recovery rate, patch-wise watermark recovery rate, correctness, and remaining portion as the cropping size increases. In specific, the remaining portion (Rem. Portion) counts the percentage of non-cropped areas within the image, which provides a reference for evaluating the process that the patch-wise watermark recovery rate drops. In general, regardless of the in-dataset or cross-dataset settings, while consistently maintaining high correctness of watermark robustness and fragility, the patch-wise watermark recovery rate closely sticks with the remaining portion as the number of cropped patches increases. At the same time, the patch-wise watermark recovery rate is always slightly greater than the remaining portion, which demonstrates the trivial over-robustness of the watermarks of FractalForensics, possibly due to strong watermark correlations between locally adjacent patches.
Table 10. Watermark recovery performance against cropping areas of different sizes on LFW.
Num. Patches Rec. Rate (bit) Rec. Rate (patch) Correctness Rem. Portion
$1\times 1$ 99.42% 98.72% 99.59% 98.44%
$2\times 2$ 97.56% 94.70% 98.93% 93.75%
$3\times 3$ 94.40% 87.83% 97.99% 85.94%
$3\times 4$ 92.45% 83.67% 97.47% 81.25%
$4\times 4$ 89.89% 78.08% 96.77% 75.00%
$4\times 5$ 87.36% 72.57% 96.05% 68.75%
$5\times 5$ 84.15% 65.65% 95.17% 60.94%
Random 95.04% 89.26% 98.12%
In Figure 12, we further visualized and compared the watermark localization performance when encountering different sizes of cropped patches. Since WaterLo (Beuve et al., 2023) provides end-to-end visualization of overlaying green and purple on the image to demonstrate real and fake areas, we mark the cropped patches as black to highlight the ground-truths. As for EditGuard (Zhang et al., 2024) and our proposed FractalForensics, besides highlighting the ground-truth cropped areas in black, we overlaid green and red for correct and incorrect localizations. In general, WaterLo is observed to have no ability to identify a small portion within the images but the entire images, with unrecognized cropped areas for the small number of cropped patches ($1\times 1$ and $2\times 2$) and recognized full images for larger numbers ($3\times 3$, $4\times 4$, and $5\times 5$). On the other hand, EditGuard identifies extra areas in addition to the correctly localized patches, while ours sometimes appears to localize fewer cropped areas than the ground-truth. Additionally, the missed patches are mostly on the contours of the cropped areas, and this is because of the consecutive ConvBlocks and SEResBlocks that strengthen correlations between adjacent patches in the model architecture of FractalForensics. However, since moderate kernel sizes are applied, no correlation is found between patches that are far apart.
(a) WaterLo (Beuve et al., 2023) against Cropping.
(b) EditGuard (Zhang et al., 2024) against Cropping.
(c) FractalForensics against Cropping.
Figure 12. Localization performance and correctness against Cropping.
Table 11. Generalizability of Z-Order fractal watermarks against common image operations for watermark recovery rate.
Dataset Fractal Identity Jpeg Gaussian Noise Gaussian Blur Median Blur Resize Average PSNR SSIM LPIPS
CelebA-HQ Hilbert 99.99% 99.97% 99.76% 99.57% 99.99% 98.03% 99.55% 35.925 0.974 0.013
Z-Order 99.99% 99.95% 99.76% 99.60% 99.87% 98.00% 99.53% 35.921 0.974 0.013
LFW Hilbert 99.95% 99.53% 99.35% 98.97% 99.50% 97.16% 99.08% 35.7405 0.9695 0.0190
Z-Order 99.94% 99.55% 99.33% 98.95% 99.49% 97.16% 99.07% 35.7460 0.9695 0.0198
Table 12. Generalizability of Z-Order fractal watermarks against Deepfake manipulations for watermark recovery rate.
Dataset Fractals SimSwap (Chen et al., 2020) InfoSwap (Gao et al., 2021) UniFace (Xu et al., 2022) E4S (Liu et al., 2023) DiffSwap (Zhao et al., 2023b) StarGAN (Choi et al., 2020) StyleMask (Bounareli et al., 2023b) HyperReenact (Bounareli et al., 2023a) Average
CelebA-HQ Hilbert 78.19% 75.12% 64.66% 69.19% 81.31% 9.92% 10.00% 9.98% 50.05%
Z-Order 78.25% 75.15% 47.64% 69.16% 81.80% 9.89% 10.09% 9.97% 47.74%
LFW Hilbert 74.87% 68.57% 51.49% 82.60% 91.48% 9.97% 10.00% 10.08% 49.88%
Z-Order 74.95% 68.52% 51.35% 82.51% 92.41% 9.90% 9.97% 10.62% 49.95%
Appendix B Generalizability of the FractalForensics
B.1. Variations of Hilbert Curve
In the main paper content, the Hilbert curve is adopted as the standard base fractal shape for demonstrations and experiments. For safety and security purposes, we ensured diversity in the candidate watermark shapes by applying variations to the standard Hilbert curve, including 3 rotations, 8 mirroring, and 3 order modifications, leading to 144 unique combinations in total. To better explain why such variations help, in Figure 13, we visualized the 14 individual variations in addition to the standard Hilbert curve. Note that although there is limited space such that we were not able to display all 144 unique fractal shapes, by looking at what is displayed, it can be easily implied that the combined variations are more complex, benefiting our goal of confidentially protecting the watermarks.
B.2. Generalization to Z-Order Curve
In this study, the semi-fragile watermarking framework is trained based on watermarks that are variations of the standard Hilbert curve. However, in practice, a user may wish to select the type of fractal shapes. Therefore, in this section, we conducted watermark generation based on another fractal shape that properly fits our watermark sizes, namely the Z-Order curve (Bader, 2000). Due to the variation capacity, we applied the 3 rotations and 8 mirroring while omitting the order modifications, leading to 36 unique variation combinations. Again, due to limited space, we visualized the individual variations along with the standard Z-Order curve in Figure 14 for better understanding.
Upon preparing watermarks based on the Z-Order curves, we tested the proposed FractalForensics in terms of watermark recovery rate regarding robustness and fragility. The results are shown in Table 11 and Table 12. For the robustness against common image operations, it can be concluded that our FractalForensics is not affected by unseen fractal shapes although the average values for Z-Order curves, 99.53% and 99.07%, are slightly lower than those for Hilbert curves at 99.55% and 99.08%, respectively. On the other hand, the Z-Order-based watermarks perform regularly against Deepfake manipulations and there are still clear gaps between robustness and fragility.
Figure 13. Individual variations of the standard Hilbert curve with respect to rotations, mirroring, and order modification.
Figure 14. Individual variations of the standard Z-Order curve with respect to rotations, mirroring, and order modification.
Appendix C Explanation on Watermark Confidentiality
Figure 15. Visualization of the bifurcation diagram for $a\in[3,4]^{\mathbb{R}}$ with $x_{0}=0.1$. Dots with the same $a$ value display the values in the encrypted sequence.
In this study, a chaotic encryption system is built to securely protect the watermarks against individuals who are not confidential. In this section, we demonstrated why the selected parameter ranges provide unpredictability and irreversibility of the one-way encryption. Regarding the logistic map
(8) $x_{i}=ax_{i-1}(1-x_{i-1}),$
with user-selected $x_{0}\in[0.1,0.9]^{\mathbb{R}}$ and $a\in[3.7,4.0)^{\mathbb{R}}$. The progression of the sequence of values is of no pattern, which can only be reproduced via the same set of parameters. In Figure 15, we plotted the bifurcation diagram of the logistic map with $x_{0}=0.1$ for the continuous range $a\in[3,4]^{\mathbb{R}}$. Considering the restriction such that $k\in[100,1000]^{\mathbb{Z}}$, we omitted $x_{i}$ for $i\in[0,99]^{\mathbb{Z}}$, and plotted 200 consecutive $x_{i}$ values for each $a$ value by setting $k=100$. It can be observed that the chaotic characteristic starts to occur after around $a=3.6$, where $a\in[3.7,4.0)^{\mathbb{R}}$ is a safe range for encryption.